OpenStudy (anonymous):

Which equation would best help solve this problem? Five added to 4 times a number is equal to 9 less than two times the number.

OpenStudy (shamim):

Let the number is x

OpenStudy (shamim):

4x+5=2x-9

OpenStudy (shamim):

X=?

OpenStudy (shamim):

-9-5=? @Nurali

OpenStudy (ccswims):

-9 - 5 = -14 not -13

OpenStudy (ccswims):

so x = -7 actually

OpenStudy (nurali):

5+4x=2x-9----(1) Subtract both side by 2x 5+4x-2x=2x-2x-9 5+2x=-9 subtract both side by 5 2x+5-5=-9-5 2x=-14 divide both side by 2 x=-7

OpenStudy (nurali):

Welcome to OpenStudy :D You may want to read the code of conduct ( http://openstudy.com/code-of-conduct ). Don't forget to click on the button called "Best Answer" to give a medal to the person who helped you :)